# Embedding ample semigroups as $(2,1,1)$-subalgebras of inverse semigroups

### v2

@@ 1. Motivation and preliminaries
Using the right ampleness of $S$ we have
$(s^{-1}s)\rho_{{}_{T}}=\rho_{s^{-1}s}=\rho_{{}_{s}}^{-1}\rho_{{}_{s}}.$
Because $\rho_{{}_{s}}^{-1}\rho_{{}_{s}}$ is identity on $Ts^{-1}s$, it follows from the definition of $\theta$ that
$(s^{-1}s)\hat{\rho}_{{}_{S}}=\left(\rho_{s^{-1}s}\right)\theta=\sigma_{s^{-1}s}$
is identity on $Ts^{-1}s\cap S$.
On the other hand $\sigma_{s^{-1}}\sigma_{s}$ is identity on $Ss$, which by the assumption is equal to $Ts^{-1}s\cap S$. Thus, we have
$(s^{-1}s)\hat{\rho}_{{}_{S}}=\left(\rho_{s^{-1}s}\right)\theta=\sigma_{s^{-1}s}=\sigma_{s^{-1}}\sigma_{s},$
implying that $(S)\hat{\rho}_{S}$ is right ample in $\mathcal{I}_{S}$. That $(S)\hat{\rho}_{S}$ is left ample in $\mathcal{I}_{S}$ follows from Theorem 2.4. ∎
Problem 2.6.
What is the situation of Theorem 2.5 with rich ample semigroups?
Similarly, using the dual notion of left $S$-invariance, one can prove that any right ample subsemigroup $S$ of an inverse semigroup $T$ can be embedded in the symmetric inverse semigroup $\mathcal{I}_{S}$ via the monomorphism,
$\hat{\lambda}_{S}:S\longrightarrow\mathcal{I}_{S},\hskip 5.0pt\text{given by }s\longmapsto{{}_{s}\sigma}={\lambda_{s}}|_{s^{-1}sT\,\cap\,S},\,\forall\,s\in S,$
where $\lambda_{s}\in\mathcal{I}_{T}$ is the partial bijection defined in Theorem 1.2. Because, in this case, $s^{-1}sT\cap S=s^{-1}sS$, we have the following analogue of Theorem 2.4.
Theorem 2.7.
Any right ample semigroup $U$ may be embedded in the symmetric inverse semigroup $\mathcal{I}_{U}$ via,
$\hat{\lambda}_{U}:U\longrightarrow\mathcal{I}_{U},\text{ given by }u\longmapsto{{}_{u}\sigma^{\prime}},$
where the partial bijection,
${{}_{u}\sigma^{\prime}}:u^{-1}uU\longrightarrow uU$
is given by $(x){{}_{u}\sigma^{\prime}}=ux$, for all $x\in u^{-1}uU$. Also, $(U){\hat{\lambda}_{U}}$ is right ample in $\mathcal{I}_{U}$.
Combining Theorems 2.4 and 2.7, we see that any ample semigroup $S$ has two isomorphic copies, viz. $(S)\hat{\rho}_{{}_{S}}$ and $(S)\hat{\lambda}_{S}$, in $\mathcal{I}_{S}$ that are, respectively, left and right ample in $\mathcal{I}_{S}$. This also gives the amalgam $\mathcal{A}=(S;\mathcal{I}_{S},\mathcal{I}_{S};\hat{\rho}_{{}_{S}},\hat{\lambda}_{S})$ (see, for instance, [6] for definition). Note that $\mathcal{A}$ is not known to be a special amalgam, because the isomorphism
$\left(\hat{\rho}_{{}_{S}}\right)^{-1}\circ\left(\hat{\lambda}_{S}\right):(S)\hat{\rho}_{{}_{S}}\longrightarrow{(S)\hat{\lambda}_{S}}$
may not extend to an automorphism of $\mathcal{I}_{S}$.
Theorem 2.8.
An ample semigroup $S$ is a $(2,1,1)$-type subalgebra of $\mathcal{I}_{S}$ if one of the following (equivalent) conditions is satisfied.
@@
    The isomorphism $\psi=(\hat{\rho}_{{}_{S}})^{-1}\circ\,\hat{\lambda}_{S}:(S)\hat{\rho}_{{}_{S}}\longrightarrow(S)\hat{\lambda}_{S}$ can be extended to a homomorphism (equivalently, isomorphism) from $V_{1}$ to $V_{2}$, that are the inverse hulls of $(S)\hat{\rho}_{{}_{S}}$ and $(S)\hat{\lambda}_{S}$, respectively.
@@
    If $x_{1}x_{2}\dots x_{n}=y_{1}y_{2}\dots y_{m}$ in $V_{1}$ then $(x_{1})\overline{\psi}(x_{2})\overline{\psi}\dots(x_{n})\overline{\psi}=(y_{1})\overline{\psi}(y_{2})\overline{\psi}\dots(y_{m})\overline{\psi}$ in $V_{2}$, where $V_{1}$ and $V_{2}$ are the inverse hulls mentoined in (1), and $\overline{\psi}$ is given by (1.6).
  3. (3)
    The amalgam $\mathcal{A}=(S;\mathcal{I}_{S},\mathcal{I}_{S};\hat{\rho}_{{}_{S}},\hat{\lambda}_{S})$ is weakly embeddable.
@@
Similar to the proof of Theorem 1.7. ∎
Remark 2.9.
The converse of Theorem 2.8 is not true. To see this, consider an arbitrary finite ample semigroup $S$ (see [7] for examples). Now, if the converse of Theorem 2.8 were true, then each of its Conditions (1) and (2) would be necessary and sufficient for $S$ to be a $(2,1,1)$-subalgebra of an inverse semigroup. This contradicts the undecidability of Problem 1.3 for finite semigroups (cf. Theorem 1.4 and Corollary 1.5) because the truth of Conditions (1) and (2) could be determined in finitely many steps.
Example 2.10.
An example is needed where a finite ample semigroup $S$ is a $(2,1,1)$-subalgebra of an inverse semigroup $T$ but is not such in $\mathcal{I}_{S}$.
Theorem 2.11.
Let $S$ be an ample semigroup such that for all $e\in E(S)$ there exists a bijection between the principal ideals $eS$ and $Se$. Then $S$ is ample in (and hence a $(2,1,1)$-subalgebra of) $\mathcal{I}_{S}$.
Proof.
If $S$ is an inverse semigroup then there is nothing to prove. So, assume that $S$ is non-inverse. Let $S$ be left ample in $T_{1}$ and right ample in $T_{2}$. Let $x$ be an arbitrary element of $S$ and $\phi_{i},1\leq i\leq 2,$ denote the embeddings of $S$ in $T_{i}$. Then
$(x)\phi_{1}((x)\phi_{1})^{-1}\in(S)\phi_{1},\text{ and }\,((x)\phi_{2})^{-1}(x)\phi_{2}\in(S)\phi_{2}.$
Let
$y=((x)\phi_{1}((x)\phi_{1})^{-1})\phi_{1}^{-1},\text{ and }\,z=(((x)\phi_{2})^{-1}(x)\phi_{2})\phi_{2}^{-1}.$
Then, clearly $y,z\in E(S)$.
Next, by Theorems 2.4 and 2.7, $\hat{S}_{1}=(S)\hat{\rho}_{{}_{S}}$ and $\hat{S}_{2}=(S)\hat{\lambda}_{S}$ are, respectively, left and right ample in $\mathcal{I}_{S}$, with
$(x)\hat{\rho}_{{}_{S}}=\sigma_{x}:Sy\longrightarrow Sx$
and
$(x)\hat{\lambda}_{S}={{}_{x}{\sigma}}:zS\longrightarrow xS.$
Furthermore, from Figure 1, one may assert that
$e_{1}=(y)\hat{\rho}_{{}_{S}}=\sigma_{y}=\sigma_{x}\sigma_{x}^{-1}:Sy\longrightarrow Sy,\hskip 10.0pt\text{ for }\,yy^{-1}=y$
and
$\begin{split}e^{\prime}_{2}=(z)\hat{\lambda}_{S}={{}_{z}\sigma}&={{}_{z}\sigma}:zS\longrightarrow zS,\hskip 40.0pt\text{ since }\,z^{-1}z=z\\ \vskip 4.0pt&={{}_{x}{\sigma}}^{-1}{{}_{x}{\sigma}}:xS\longrightarrow xS\end{split}$
Let us also observe that
$(e^{\prime}_{2})\psi^{-1}=(z)\hat{\rho}_{{}_{S}}=\sigma_{z}:Sz\longrightarrow Sz$
is an idempotent in $(S)\hat{\rho}_{{}_{S}}$, and
$(e_{1})\psi=(y){{}_{S}\hat{\theta}}={{}_{y}\sigma}:yS\longrightarrow yS$
belongs to $E((S)\hat{\lambda}_{S})$. Let
$e^{\prime}_{1}=\sigma^{-1}_{x}\sigma_{x}:Sx\longrightarrow Sx,\;\;\text{(which may not belong to }(S)\hat{\rho}_{{}_{S}}),$
and
$e_{2}={{}_{x}\sigma}{{}_{x}\sigma}^{-1}:zS\longrightarrow zS,\;\;\text{(which may not be a member of }(S)\hat{\lambda}_{S}).$
We may now calculate
$\begin{split}\sigma_{x}\left((e^{\prime}_{2})\psi^{-1}\right)&=\sigma_{x}\left(({{}_{x}\sigma}^{-1}{{}_{x}\sigma})\psi^{-1}\right)\\ &=\left(({{}_{x}\sigma})\psi^{-1}\right)\left(({{}_{x}\sigma}^{-1}{{}_{x}\sigma})\psi^{-1}\right)\\ &=({{}_{x}\sigma}{{}_{x}\sigma}^{-1}{{}_{x}\sigma})(\psi)^{-1}\\ &=({{}_{x}\sigma})(\psi)^{-1}\\ &=\sigma_{x}.\end{split}$
Because $\sigma_{x}^{-1}\sigma_{x}$ is the minimal idempotent such that $\sigma_{x}(\sigma^{-1}_{x}\sigma_{x})=\sigma_{x}$, we must have
$\sigma^{-1}_{x}\sigma_{x}\leq(e^{\prime}_{2})\psi^{-1}=({{}_{x}\sigma}^{-1}{{}_{x}\sigma})\psi^{-1}=({{}_{z}\sigma})\psi^{-1}=\sigma_{z}.$ (2.6)
By a similar token, we also obtain
${{}_{x}\sigma}{{}_{x}\sigma^{-1}}\leq\,{{}_{y}\sigma}.$ (2.7)
The aim is to prove that $Dom(\sigma^{-1}_{x}\sigma_{x})=Dom(\sigma_{z})$ and $Dom({{}_{x}\sigma}{{}_{x}\sigma^{-1}})=Dom({{}_{y}\sigma})$. To this end, we first note, using inequalities (2.6) and (2.7), respectively, that
$Sx=Dom(\sigma^{-1}_{x}\sigma_{x})\subseteq Dom(\sigma_{z})=Sz$
and
$zS=Dom({{}_{x}\sigma}{{}_{x}\sigma^{-1}})\subseteq Dom({{}_{y}\sigma})=yS.$
Hence, we have, by the hypothesis,
$Sx\subseteq Sz=zS\subseteq yS=Sy.$
Now, because $S$ is finite and $\sigma_{x}:Sy\longrightarrow Sx$ is a bijection, we get
$Sx=Sz=zS=yS=Sy.$
Recalling the definition of the natural partial order, it follows from (2.6) and (2.7) that $e^{\prime}_{1}\in(S)\hat{\rho}_{{}_{S}}$ and $e_{2}\in(S)\hat{\lambda}_{S}$. This completes the proof. ∎
The following corollary is straightforward.
@@
Problem 2.13.
What are the inverse semigroups such that the ample semigroups considered in the above theorem and its corollary are their full semigroups (up to isomorphism)?
@@ 3. Examples concerning left (right) ample semigroups
A semigroup $S$ without zero is called simple if it has no proper (two-sided) ideals (recall that groups are precisely the semigroups with no proper one-sided ideals). We say that $S$ is completely simple if it is simple and contains an idempotent which is minimal within the set $E(S)$. If a semigroup $S$ has a zero element 0, then clearly $0\leq e$ for all $e\in E(S)$. In this case, an idempotent $f\in E(S)$ is called primitive if $e\leq f$ implies that $e\in\{0,f\}$ for all $e\in E(S)$. A semigroup $S$ with zero is called 0-simple if $S^{2}\neq{0}$ and 0{0} is the only proper ideal of $S$. (The first condition only serves to exclude the two-element null semigroup, which makes the relevant structure theory somewhat clean.) A semigroup $S$ with zero is said to be completely 0-simple if it is 0-simple and contains a primitive idempotent. Completely (simple) 0-simple semigroups form an important class of (regular) semigroups. They are characterized by the celebrated Rees representation theorems (cf. [5], Theorems 3.2.3 and 3.3.1). In this article we shall consider completely 0-simple inverse semigroups that are known as Brandt semigroups. (Completely simple inverse semigroups are in fact groups.) For the convenience of the reader, we are reproducing below the Rees representation theorem for completely 0-simple semigroups, as well as its version for Brandt semigroups. For further background on Rees matrix semigroups the reader may refer to [2], Chapter 3.
@@
The set $(I\times G\times\Lambda)\cup\{0\}$ together with the above binary operation is called the $I\times\Lambda$ Rees matrix semigroup over the 0-group $G^{0}$ with sandwich matrix $P$. We denote it by $\mathcal{M}^{0}(G^{0},I,\Lambda,P)$.
@@
Since the sandwich matrix $\Delta$ will always be obvious, from now on, the Brandt semigroup $B=\mathcal{M}^{0}(G^{0},I,I,\Delta)$ will be denoted by its universe $(I\times G\times I)\cup\{0\}$. Note that the inverse of $(i,g,j)\in B$ is $(j,g^{-1},i)$, where $g^{-1}$ is the inverse of $g$ in $G$. Also, the non-zero idempotents of $B$ are of the form $(i,1,i),\,i\in I$, where $1$ is the identity of $G$.
It is an easy exercise to verify that a semigroup with zero that has no proper non-zero one sided ideal is a zero group (and hence an inverse semigroup). This implies that commutative 0-simple semigroups, equivalently, the commutative Brandt semigroups, are zero (Abelian) groups. It is also a routine to verify that a Brandt semigroup is a zero group if and only if $I$ is a singleton. In this section we shall consider left (right) ample subsemigroups of a Brandt semigroup, such that the latter is not a zero group. So, we shall only be dealing with non-commutative Brandt semigroups. This also means that the set $I$ will contain more than one element.
@@
    $(S)\hat{\rho}_{{}_{S}}$ is left ample in $\mathcal{I}_{S}$.
@@
    For every inverse oversemigroup $T$ of $S$, either $\hat{\lambda}_{{}_{S}}$, given by Theorem 2.7, is not a monomorphism or $(S)\hat{\lambda}_{{}_{S}}$ is not right ample in $\mathcal{I}_{S}$.
@@
($\implies$) Let $S$ be strict left ample. Then, by Theorem 2.4, $(S)\hat{\rho}_{{}_{S}}$ is left ample in $\mathcal{I}_{S}$. Also, if $\hat{\lambda}_{S}$ is an isomorohism and $(S)\hat{\lambda}_{S}$ is right ample in $\mathcal{I}_{S}$ then $S$ is indeed right ample, a contradiction.
@@
Let us arbitrarily fix an element $e=(\lambda,1,\lambda)\in E(B)\smallsetminus\{0\}$. We show that $eB$ is a strict left ample subsemigroup of $B$. (One may prove analogously that $Be$ is strict right ample in $B$.) Let $u$ be an arbitrary element of $eB$. If $u=0$, then $uu^{-1}=0\in eB$ and there is nothing to prove. So, assume that $u\in eB\smallsetminus\{0\}$. Then, we have
@@
Now, one can easily verify that $uu^{-1}=(\lambda,1,\lambda)$, where $u^{-1}$ denotes the inverse of $u$ in $B$. Clearly, $(\lambda,1,\lambda)$ is an element of $eB$. Hence, $eB$ is left ample (in $B$).
Next, we use Proposition 3.4 to prove that $S=eB$ is not right ample. The aim is to show that $\hat{\lambda}_{S}$ is not a monomorphism. Consider, for this purpose, $u=(\lambda,h,l)\in eB\smallsetminus\{0\}$. Let $u^{-1}$ denote the inverse of $u$ in some (up to isomorphism) inverse oversemigroup $T$ of $eB$, such that $u^{-1}u\in eB$ (if no such inverse semigroup exists then we are done). Then, after effectuating necessary identifications, one can easily verify from $uu^{-1}u=u\neq 0$ that $u^{-1}u=(l,1,l)$. Because $I$ is not a singleton, we can choose $u$ such that $l\neq\lambda$. But then,
$\begin{split}\hat{\lambda}_{u}&=(l,1,l)(eB)\longrightarrow(\lambda,h,l)(eB)\\ &=(l,1,l)(\lambda,1,\lambda)B\longrightarrow(\lambda,h,l)(\lambda,1,\lambda)B\\ &=0B\longrightarrow 0B\\ &=\hat{\lambda}_{0}\end{split},$
implying the $\hat{\lambda}_{S}$ is not a monomorphism. ∎
The following example provides a left ample semigroup of a Brandt semigroup that is not contained in any of the non-zero idempotent generated ideals.
@@
Let $\mathbb{Q}^{+}$ be the multiplicative group of positive rationals, $\mathbb{N}_{1}$ the multiplicative semigroup of natural numbers starting from 1 and $I=\{1,2,3\}$. Consider the Brandt semigroup $B=(I\times\mathbb{Q}^{+}\times I)\cup\{0\}$ and its subsemigroup $S=(\{1,2\}\times\mathbb{N}_{1}\times I)\cup\{0\}$. Subsemigroup $S$ is left ample, because for any $n\in\mathbb{N}_{1}$, $j\in\{1,2\}$ and $i\in I$, we have $(j,n,i)(i,n^{-1},j)=(j,1,j)\in S$, but not right ample, because $(3,n^{-1},j)(j,n,3)=(3,1,3)\not\in S$. The non-zero idempotent generated ideals are $(i,1,i)B=(\{i\}\times\mathbb{Q}^{+}\times I)\cup\{0\}$. Clearly $S$ is not contained in any of them.
Remark 3.7.
The construction used in the above example works for any group (including finite ones). Let $G$ be a group and $I=\{1,2,3\}$. Consider the the Brandt semigroup $B=(I\times G\times I)\cup\{0\}$ and the subsemigroup $S=(\{1,2\}\times G\times I)\cup\{0\}$. (Actually we may take $S=(\{1,2\}\times H\times I)\cup\{0\}$, where $H$ is any submonoid of $G$.) Now we have, for any $g\in G$, $j\in\{1,2\}$ and $i\in I$, $(j,g,i)(i,g^{-1},j)=(j,1,j)\in S$, but $(3,g^{-1},j)(j,g,3)=(3,1,3)\not\in S$. Hence $S$ is strict left ample. On the other hand $S\not\subset(\{i\}\times G\times I)\cup\{0\}$ for any $i\in I$, but all principle ideals are of the form $(\{i\}\times G\times I)\cup\{0\}$. The reason for the previous example to work is that the strict ample subsemigroup (also a right ideal) is generated by two idempotents $(1,1,1)$ and $(2,1,2)$.
Problem 3.8.
Can we prove that Theorem 3.5 does not apply to the above example and remark?
Let $B=(I\times G\times I)\cup\{0\}$ be a Brandt semigroup. Then clearly every subsemigroup of $B$ with zero is of the form $S=(I^{\prime}\times H\times I^{\prime\prime})\cup\{0\}$ where $I^{\prime}$ and $I^{\prime\prime}$ are subsets of $I$ and $H$ is a subsemigroup of the group $G$; a subsemigroup without zero is of the form $I^{\prime}\times H\times I^{\prime\prime}$.

### v3

@@ 2. Answering the embedding problem for some finite ample semigroups
If $S$ is an inverse semigroup then there is nothing to prove. So, assume that $S$ is a finite non-inverse ample semigroup satisfying the condition of the theorem.
Let $\phi_{i}:S\longrightarrow T_{i},\,1\leq i\leq 2$, be monomorphisms such that $S_{1}=S\phi_{1}$ is left ample in $T_{1}$ and $S_{2}=S\phi_{2}$ is right ample in $T_{2}$. Also, for each $i\in\{1,2\}$, let $\phi^{\prime}_{i}:S\longrightarrow S_{i}$ be the isomorphisms defined by $s\longmapsto x\phi_{i}$. In a way similar to (1.4), define for all $x\in S$, $x_{1}x_{1}^{-1},x_{2}^{-1}x_{2}\in E(S)$ by
$\begin{split}x_{1}x_{1}^{-1}&:=[(x\phi_{1})(x\phi_{1})^{-1}]\phi_{1}^{\prime-1},\vskip 4.0pt\\ x_{2}^{-1}x_{2}&:=[(x\phi_{2})^{-1}(x\phi_{2})]\phi_{2}^{\prime-1}.\end{split}$
Also, denote $S\hat{\rho}$ and $S\hat{\lambda}$ by $\hat{S_{1}}$ and $\hat{S_{2}}$, respectively, where $\hat{\rho}$ is the monomorphism given by Theorems 2.6 and $\hat{\lambda}$ is the anti-monomorphism given by Theorem 2.7. Then there exists an anti-isomorphism $\psi:\hat{S_{1}}\longrightarrow\hat{S_{2}}$ such that $\hat{\rho}\circ\psi=\hat{\lambda}$. Note that $f=\sigma\circ\chi:S_{1}\longrightarrow\mathcal{I}_{S}$ is a monomorphism (refer to Figure 1). Also, there exists an anti-monomorphism $g:S_{2}\longrightarrow\mathcal{I}_{S}$ yielded from Theorem 2.7. (All these maps are shown in Figure 2.)
Now, consider the following elements of $\hat{S_{1}}$:
$\begin{split}\hat{\rho}_{x}&:Sx_{1}x_{1}^{-1}\longrightarrow Sx,\\ e_{1}=\hat{\rho}_{x_{1}x_{1}^{-1}}&:Sx_{1}x_{1}^{-1}\longrightarrow Sx_{1}x_{1}^{-1}\\ =\hskip 1.5pt\hat{\rho}_{x}\hat{\rho}^{-1}_{x}&:Sx_{1}x_{1}^{-1}\longrightarrow Sx_{1}x_{1}^{-1}.\end{split}$
Analogously, the following elements belong to $\hat{S_{2}}$:
$\begin{split}\hat{\lambda}_{x}&:x_{2}^{-1}x_{2}S\longrightarrow xS,\\ e_{2}=\;\hat{\lambda}_{x_{2}^{-1}x_{2}}&:x_{2}^{-1}x_{2}S\longrightarrow x_{2}^{-1}x_{2}S\\ =\hskip 3.5pt\hat{\lambda}_{x}\hat{\lambda}^{-1}_{x}&:x_{2}^{-1}x_{2}S\longrightarrow x_{2}^{-1}x_{2}S.\end{split}$
As $\psi(\hat{\rho}_{x})=\hat{\lambda}_{x}$, we also note that
$\begin{split}(e_{2})\psi^{-1}=\hat{\rho}_{x_{2}^{-1}x_{2}}&:S{x_{2}^{-1}x_{2}}\longrightarrow S{x_{2}^{-1}x_{2}}\in\hat{S}_{1},\\ (e_{1})\psi=\hat{\lambda}_{x_{1}x_{1}^{-1}}&:x_{1}x_{1}^{-1}S\longrightarrow x_{1}x_{1}^{-1}S\in\hat{S}_{2}.\end{split}$ (2.5)
Let us now carry out the following calculation:
$\begin{split}(\hat{\rho}_{x})(e_{2}\psi^{-1})&=(\hat{\rho}_{x})\left((\hat{\lambda}_{x}\hat{\lambda}^{-1}_{x})\psi^{-1}\right)\\ &=(\hat{\lambda}_{x})\psi^{-1}\left((\hat{\lambda}_{x}\hat{\lambda}^{-1}_{x})\psi^{-1}\right)\\ &=(\hat{\lambda}_{x}\hat{\lambda}^{-1}_{x}\hat{\lambda}_{x})\psi^{-1}\\ &=(\hat{\lambda}_{x})\psi^{-1}\\ &=\hat{\rho}_{x}\end{split}$
Because $\hat{\rho}_{x}^{-1}\hat{\rho}_{x}\in E(\mathcal{I}_{S})$ is the minimal idempotent such that $\hat{\rho}_{x}(\hat{\rho}^{-1}_{x}\hat{\rho}_{x})=\hat{\rho}_{x}$, we must have
[…]
By a similar token, we also obtain
${\hat{\lambda}_{x}^{-1}}{\hat{\lambda}_{x}}\leq\,(e_{1})\psi={\hat{\lambda}_{x_{1}x^{-1}_{1}}}.$ (2.7)
The proof will be accomplished if show that $\hat{\rho}^{-1}_{x}\hat{\rho}_{x}\in\hat{S}_{1}$, wherefore it suffices to show that
$Dom(\hat{\rho}^{-1}_{x}\hat{\rho}_{x})=Dom(\hat{\rho}_{x^{-1}_{2}x_{2}})\hskip 8.0pt(=Dom(e_{2})\psi^{-1}).$
Since $S$ is a finite semigroup and by (2.6) we have $Dom(\hat{\rho}^{-1}_{x}\hat{\rho}_{x})\subseteq Dom(\hat{\rho}_{x^{-1}_{2}x_{2}})$, the aim will be achieved if there exists a bijection between $Dom(\hat{\rho}^{-1}_{x}\hat{\rho}_{x})$ and $Dom(\hat{\rho}_{x^{-1}_{2}x_{2}})$. Let $\mid X\mid$ denote the cardinality of $X$. Then the aim is to show that
$\mid Dom(\hat{\rho}^{-1}_{x}\hat{\rho}_{x})\mid\,=\,\mid Dom(\hat{\rho}_{x^{-1}_{2}x_{2}})\mid.$
Using (2.6) and (2.7) we first obtain,
$\mid Sx\mid\,=\,\mid Dom(\hat{\rho}^{-1}_{x}\hat{\rho}_{x})\mid\,\leq\,\mid Dom(\hat{\rho}_{x_{2}^{-1}x_{2}})\mid\,=\,\mid Sx_{2}^{-1}x_{2}\mid,$
and
$\mid xS\mid\,=\,\mid Dom({\hat{\lambda}_{x}^{-1}}{\hat{\lambda}_{x}})\mid\,\leq\,\mid Dom({\hat{\lambda}_{x_{1}x_{1}^{-1}}})\mid\,=\,\mid x_{1}x_{1}^{-1}S\mid.$
Now, because $\mid x_{2}^{-1}x_{2}S\mid\,=\,\mid xS\mid\,$ and $\mid Sx_{1}x_{1}^{-1}\mid\,=\,\mid Sx\mid$, we obtain, by the hypothesis:
$\mid Sx\mid\,\leq\,\mid Sx_{2}^{-1}x_{2}\mid\,=\,\mid x_{2}^{-1}x_{2}S\mid\,=\,\mid xS\mid\,\leq\,\mid x_{1}x_{1}^{-1}S\mid\,=\,\mid Sx_{1}x_{1}^{-1}\mid\,=\,\mid Sx\mid.$ (2.8)
This implies that
$\mid Dom(\hat{\rho}^{-1}_{x}\hat{\rho}_{x})\mid\,=\,\mid Sx\mid\,=\,\mid Sx^{-1}_{2}x_{2}\mid\,=\,\mid Dom(\hat{\rho}_{x^{-1}_{2}x_{2}})\mid.$
and the proof is complete. ∎
Corollary 2.9.
If $S$ is an ample semigroup then the following conditions are equivalent.
@@
    For every $e\in E(S)$ There exists a bijection between the principal ideals $eS$ and $Se$.
@@
    For all $x\in S$ there exists a bijectoon between $xS$ and $Sx$.
@@
[…]
The class of all finite semigroups with central idempotents froms a pseudovariety, denoted by $\mathbf{ZE}$. Denoting by $\mathbf{G}$ and $\mathbf{Com}$ the pseudovarieties of groups and commutative semigroups, we have [2]:
$\mathbf{ZE}=\mathbf{G}\vee\mathbf{Com}.$
If we denote the class (in fact quasivariety) of ample semigroups by $\mathbf{A}$ then the following corollary shows that every member of $\mathbf{A}\wedge\mathbf{ZE}$ is embeddable in an inverse semigroup as $(2,1,1)$-subalgebra.
@@
Proof.
Straightforward. ∎
@@ 3. Examples concerning left (right) ample semigroups
[…]
@@
The set $(I\times G\times\Lambda)\cup\{0\}$ together with the above binary operation is called the $I\times\Lambda$ Rees matrix semigroup over the 0-group $G^{0}$ with sandwich matrix $P$. We denote it by $\mathcal{M}^{0}(G^{0},I,\Lambda,P)$. For further background on Rees matrix semigroups the reader may refer to [10], Chapter 3.
@@
Since the sandwich matrix $\Delta$ will always be obvious, the Brandt semigroup $\mathcal{M}^{0}(G^{0},I,I,\Delta)$ will, from now on, be denoted by its universe $(I\times G\times I)\cup\{0\}=B$. Note that the inverse of $(i,g,j)\in B$ is $(j,g^{-1},i)$, where $g^{-1}$ is the inverse of $g$ in $G$. Also, the non-zero idempotents of $B$ are of the form $(i,1,i),\,i\in I$, where $1$ is the identity of $G$.
It is an easy exercise to verify that a semigroup with zero that has no proper non-zero one sided ideal is a zero group (and hence an inverse semigroup). This implies that commutative 0-simple semigroups, equivalently, the commutative Brandt semigroups, are zero (Abelian) groups. It is also a routine to verify that a Brandt semigroup is a zero group if and only if $I$ is a singleton. We shall consider left (right) ample subsemigroups of a Brandt semigroup, such that the latter is not a zero group. So, we shall only be dealing with non-commutative Brandt semigroups such that the set $I$ contains more than one element.
@@
    $(S)\hat{\rho}=\hat{S}_{1}$ is left ample in $\mathcal{I}_{S}$, refer to Theorem 2.6 and Figure 2.
@@
    Either $\hat{\lambda}$ is not an anti-monomorphism or $(S)\hat{\lambda}=\hat{S}_{2}$ is not left ample in $\mathcal{I}_{S}$, cf. Theorem 2.7 and Figure 2.
Dual conditions hold for strict right ample semigroups.
@@
($\implies$) Let $S$ be strict left ample. Then, by Theorem 2.6, $\hat{S}_{1}$ is left ample in $\mathcal{I}_{S}$. To prove condition (2), assume on the contrary that $\hat{\lambda}$ is an anti-monomorohism and $\hat{S}_{2}$ is left ample in $\mathcal{I}_{S}$. Consider
$\hat{S}^{-1}_{2}=\{x^{-1}:x\in\hat{S}_{2}\},$
a subsemigroup of $\mathcal{I}_{S}$. Then, $h:\hat{S}_{2}\longrightarrow\hat{S}^{-1}_{2}$, given by $x\mapsto x^{-1}$, is clearly an anti-isomorphism, and $\hat{S}^{-1}_{2}$ is right ample in $\mathcal{I}_{S}$. But then $\hat{\lambda}\circ h:S\longrightarrow\mathcal{I}_{S}$ is a monomorphism, implying that $S$ is right ample, a contradiction.
@@
Let us arbitrarily fix an element $e=(\lambda,1,\lambda)\in E(B)\smallsetminus\{0\}$. We shall only show that $eB$ is a strict left ample subsemigroup of $B$ (the statement about $Be$ can be proved by a dual argument). Let $u$ be an arbitrary element of $eB$. If $u=0$, then $uu^{-1}=0\in eB$ and there is nothing to prove. So, assume that $u\in eB\smallsetminus\{0\}$. Then, we have
@@
Now, one can easily verify that $uu^{-1}=(\lambda,1,\lambda)$, where $u^{-1}$ denotes the inverse of $u$ in $B$. Clearly, $(\lambda,1,\lambda)$ is an element of $eB$. Hence, $eB$ is left ample in $B$.
Next, we use Remark 3.4 to prove that $S=eB$ is not right ample. The aim is to show that $\hat{\lambda}:S\longrightarrow\mathcal{I}_{S}$ is not an anti-monomorphism. Consider, for this purpose, an element $u=(\lambda,h,l)\in eB\smallsetminus\{0\}$. Let $u^{-1}$ denote the inverse of $u$ in some inverse oversemigroup $T$ of $eB$, such that $u^{-1}u\in eB$ (if no such inverse semigroup exists then we are done). Then, after effectuating necessary identifications, one can easily verify from $uu^{-1}u=u\neq 0$ that $u^{-1}u=(l,1,l)$. Because $I$ is not a singleton, we can choose $u$ such that $l\neq\lambda$. But then,
$\hat{\lambda}_{u}:(l,1,l)(eB)\longrightarrow(\lambda,h,l)(eB)$
coincides with
$\hat{\lambda}_{0}:0B\longrightarrow 0B,$
because
$\begin{split}(l,1,l)(eB)=(l,1,l)(\lambda,1,\lambda)B=0B,\vskip 4.0pt\\ (\lambda,h,l)eB=(\lambda,h,l)(\lambda,1,\lambda)B=0B,\end{split}$
implying the $\hat{\lambda}$ is not an anti-monomorphism. This implies, by Proposition 3.4, that $S=eB$ is not right ample. ∎
The following example provides a class of strict left ample subsemigroups of Brandt semigroups that are not contained in any non-zero idempotent generated ideals.
@@
Consider a Brandt semigroup $B=(I\times G\times I)\cup\{0\}$, where $G$ is group and $I=\{1,2,3\}$. Let $S=(\{1,2\}\times H\times I)\cup\{0\}$, where $H$ is a submonoid of $G$. Then we have, for all $g\in G$, $j\in\{1,2\}$ and $i\in I$, $(j,g,i)(i,g^{-1},j)=(j,1,j)\in S$. So, identifying $S$ with $S\hat{\rho}$, we see that Condition (1) of Remark 3.4 is satisfied. To prove Condition (2), we identify $S$ with $S\hat{\lambda}$ and notice that $(3,g^{-1},j)(j,g,3)=(3,1,3)\not\in S$. Hence, by Remark 3.4, $S$ is strict left ample. At the same time, all principal ideals of $B$ are of the form $(\{i\}\times G\times I)\cup\{0\}$ and $S\not\subset(\{i\}\times G\times I)\cup\{0\}$ for any $i\in I$.
Let $B=(I\times G\times I)\cup\{0\}$ be a Brandt semigroup. Then clearly every subsemigroup of $B$ with zero is of the form $S=(I^{\prime}\times H\times I^{\prime\prime})\cup\{0\}$ where $I^{\prime}$ and $I^{\prime\prime}$ are subsets of $I$ and $H$ is a subsemigroup of the group $G$. Obviously, a subsemigroup of $B$ without zero is of the form $I^{\prime}\times H\times I^{\prime\prime}$.